J TAX (Dolars per Dvo As the tax per DVD increases, deadweight loss

4. The Laffer curve Governments often place so-called sin taxes on goods or services such as cigarettes, alcohol, and pornography. These kinds of taxes a palatable to voters than income taxes ular with politicians because they are usually more re pop To understand the effect of such a tax, consider the monthly market for adult DVDs, which is shown on the following graph. Use the graph input tool to help you answer the following questions. You will not be graded on any changes you make to this graph. Note: Once you enter a value in a white field, the graph and any corresponding amounts in each grey field will change accordingly Graph Input Tool Market for Adult DVDs Supply 80 (DVDs) Demand Price Dollars per 24.00 Supply Price 16.00 Dollars per VD) Tax Wedge Dollars per VD) 8.00 Suppose the government imposes an $8-per-DVD tax on suppliers. At this tax amount, the equilibrium quantity of adult DVDs is DVDs, and the government collects S in tax revenue Now caiculate the government's tax revenue if it sets a tax of $0, $8, $16, $20, $24, $32, or $40 per DVD. (Hint: To find the equilibrium quantity after the tax, adjust the Quantity"field until the Tax Wedge equais the value of the per-unit tax.) Using the data you generate, piot a Laffer curve by using the green points (triangle symbol) to plot total tax revenue at each of those tax levels. Note: Plot your points in the order in which you would like them connected. Line segments will connect the points automatically
980 TAX (Dlas per DND Suppose the government is currently imposing a $24-per-DVD tax on adult DVDs. True or False: The government can raise its tax revenue by decreasing the per-unit tax on adult DVDs. False Consider the deadweight loss generated in each of the following cases: no tax, a tax of $16 per DVD, and a tax of $32 per DVD. On the following graph, use the black curve (plus symbols) to ilustrate the deadweight loss in these cases. (Hint: Remember that the area of a triangle is equal to 1xBare x Height. In the case of a deadweight loss triangle found on the graph input tool, the base is the amount of the tax and the height is the reduction in quantity caused by the tax.)
"J TAX (Dolars per Dvo As the tax per DVD increases, deadweight loss
Answer #1

Solution:

At the tax of $8 per unit, it can be seen from the given graph that equilibrium quantity is 80 DVDs. The tax revenue collected by government = tax amount*equilibrium quantity

Tax revenue = 8*80 = $640

By filling in the values of different tax rates, we can find the equilibrium level of quantity (appearing in the shown blank):

At tax rate of $0, that is not tax and no government intervention, free market equilibrium is achieved. Thus, equilibrium quantity = 100 DVDs; so tax revenue = 0*100 = $0

At tax rate of $16, equilibrium quantity is 60 DVDs, so tax revenue = 16*60 = $960

At tax rate of $20, equilibrium quantity is 50 DVDs, so TR = 20*50 = $1000

At tax rate of $24, equilibrium quantity is 40 DVDs, so TR = 24*40 = $960

At tax rate of $32, equilibrium quantity is 20 DVDs, so TR = 32*20 = $640

At tax rate of $40, equilibrium quantity is 0 DVDs, so TR = 40*0 = $0

Plotting the points: (0, 0), (8, 640), (16, 960), (20, 1000), (24, 960), (32, 640) and (40, 0), we obtain the Laffer curve:

TAX REVENUE 2500 2000 E 1500 1000 500 0 5 10 15 20 25 30354045 Tax

At the per unit tax rate of $24, the tax revenue collected is $960. Clearly, from the graph above, this point lies on the downward sloping side of the Laffer curve, thus by reducing the tax rate, the tax revenue collected can be increased. So, the given statement is TRUE.

Dead weight loss is simply the welfare or efficiency loss caused due to taxation:

So, as the definition of dead weight loss suggests, with no tax there is no dead weight loss generated (full efficiency or maximum social welfare possible is generated). So, DWL = $0

At tax rate of $16, DWL = (1/2)*(tax rate)*(optimal quantity - equilibrium quantity)

So, DWL = (1/2)*16*(100 - 60) = $320

At tax rate of $32, DWL = (1/2)*32*(100 - 20) = $1280

So, the following graph is obtained:

Dead weightloss 1400 1200 1000 E 800 600 400 200 10 15 20 25 30 35 ax

Notice that this curve is upward sloping, indicating that as the per unit tax rate increases, dead weight loss increases.